[Brahms-l] Shift report 20040219 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Thu Feb 19 2004 - 23:39:20 EST
------------------------------------------------------------
Date: Thu Feb 19 23:39:19 2004, Shift: 16:00-24:00
Supervisor: BHS, with: PJ
Log book#15 pages 62 - 62

Shift Summary:
------------------------------------------------------------
1: Machine delevopment and a short access for the first hour of the shift, then one really nice, long store. Initial intensities close to 50/50 :-) One QLI as they were refilling at 11.15. Also, RHIC fill pattern has been changed so BRAHMS will (for the next store at least) only recieve 52 bunches instead of 56.

2: Took four data runs (10425-10428) at 4deg 1/2A, 45 6A. Trigger count:
Tr.2: 550k
Tr.3: 860k

3: No detector problems. We need a new hole-puncher though - the old one has finally started falling apart and will now only make two holes (even after attempts at repair...)
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Thu Feb 19 23:39:27 2004

This archive was generated by hypermail 2.1.8 : Thu Feb 19 2004 - 23:39:49 EST